The length of a rectangle is three less than two times the width. If the perimeter is 42, find the length and width. (Equation and solve)
Answer by cleomenius(959)

l = 2x - 3
w = x
2(2x - 3) + 2x = 42
4x - 6 + 2x = 42
6x = 48
x = 8 the width.
2(8) -3 = 13, the length.
16 + 26 = 42.
Cleomenius.